OpenStudy (whynothelp1137):

The solution set for all real numbers x such that 2x + 3 < 19 and 2x − 3 ≤ -7 is (, ). Note: If needed, type “-infinity” for -∞ and “infinity” for ∞.

OpenStudy (whynothelp1137):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

first just focus on solving 2x + 3 < 19 (ignore the other inequality)

OpenStudy (whynothelp1137):

Alright.

OpenStudy (whynothelp1137):

When x= anything negitive wouldn't it be less than 19?

jimthompson5910 (jim_thompson5910):

how would you solve something like 2x+3 = 19 ?

OpenStudy (whynothelp1137):

I'd solve it by pluging in numbers for X?

jimthompson5910 (jim_thompson5910):

so a guess and check method?

OpenStudy (whynothelp1137):

Practically, yes.

jimthompson5910 (jim_thompson5910):

that might take a while, why not subtract 3 from both sides 2x+3 = 19 2x+3-3 = 19-3 2x+0 = 16 2x = 16 then divide both sides by 2 2x = 16 2x/2 = 16/2 x = 8

jimthompson5910 (jim_thompson5910):

To solve equations, you follow PEMDAS in reverse and undo everything to isolate the variable you want

jimthompson5910 (jim_thompson5910):

when I subtracted 3 from both sides, I did that to undo the +3 when I divided both sides by 2, I undid the multiplication of 2

OpenStudy (whynothelp1137):

I seen that, So then X would just equal 8? Nothing less?

jimthompson5910 (jim_thompson5910):

well that's the equation form

jimthompson5910 (jim_thompson5910):

we're dealing with inequalities though

jimthompson5910 (jim_thompson5910):

use the same steps to solve 2x+3 < 19 2x+3 < 19 2x+3-3 < 19-3 2x < 16 2x/2 < 16/2 x < 8 make sense?

OpenStudy (whynothelp1137):

It does. Since pluging 8 in as x would cause it to be 19 we need less than 19.

jimthompson5910 (jim_thompson5910):

solve 2x − 3 ≤ -7 and tell me what you get

OpenStudy (whynothelp1137):

X=-2

jimthompson5910 (jim_thompson5910):

it's not an equation though

jimthompson5910 (jim_thompson5910):

you should get \(\Large x \le -2\)

OpenStudy (whynothelp1137):

Ah. My bad.

jimthompson5910 (jim_thompson5910):

so solving the two inequalities 2x + 3 < 19 and 2x − 3 ≤ -7 leads to x < 8 and x ≤ -2

jimthompson5910 (jim_thompson5910):

the keyword "and" is very important here

jimthompson5910 (jim_thompson5910):

think of a number that is BOTH smaller than 8 AND less than or equal to -2 what number do you come up with?

OpenStudy (whynothelp1137):

Well, that could be anything beyond -2 could it not?

jimthompson5910 (jim_thompson5910):

Let's use a number line

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

the graph of x < 8 would look like this  it has an open hole at 8 and the shading is to the left